# “Add more config detail”: A Taxonomy of Installation Instruction Changes

### v1

@@ 4.3 Post-installation Instruction
In addition, OSS projects sometimes contain multiple runnable modules, each serving a specific aspect of the project, such as server/worker modules in distributed system projects and training/inference modules in machine learning projects. Properly documenting each component’s execution instruction is essential for users to make use of the software product. Within our analysed data, we identified in Example 15 that additional execution commands were added for other blocks of functions. This could happen to remedy incomplete documentation issues [5] or to reflect newly added features in the system.
Instruction updates for software execution are also common in our analysed data. Changes are found to be frequently appearing in changing file names or file path of the executable, changing default parameter settings for the instruction. These changes are highly intertwined with code base changes, giving rise to motivation of automated documentation tools to perform updates.
@@
Summary of post-installation instruction results: The most dominant concept in this category is project running instruction. Changes in this concept are found to be related to updating parameter options and file paths to the executables. Other concepts are less frequently encountered but they all constitue the installation-related software development tasks.
4.4 Help Information
@@
Tutorial Updates (70). Limited in space and purpose, README files typically do not offer comprehensive usage instructions, such as API documentation. We observed that most updates within this concept involve providing code demos for basic usage of a project. Most documentation changes are trivial, merely made to align with the source code modifications. In Example 17, updates are made within the code demos, altering all naming paradigms for the methods of using the model.
@@
Summary of help information results: Tutorials cover basic project usage with code demos. Explanation information serves as complementary materials to instructions, providing guides when instructions are optional or not common. Notes and troubleshooting provides additional information, and are dedicated to providing guidance towards common errors for the software project.
4.5 Document Presentation
Different from the previous sections, document presentation updates involve less content modification. Instead, the emphasis is on organising the documentation to enhance visual presentation and reading comprehension. In total, 477 of 1,168 data points are associated with this category. In this category, we synthesised five concepts, which are discussed as follows:
@@ 4.5 Document Presentation
Content Formatting and Cleaning (119). README content formatting involves using Markdown syntax to either highlight or organise the content. Various information is commonly formatted, including large code blocks in Example 24, inline code blocks in Example 25, and URLs in Example 26. All of these formatting operations contribute to enhance the presentation of the documentation, making the README files easier to follow. Specifically, the formatting on code blocks enables better highlighting on code components, while the formatting on URL facilitates enhanced accessibility and clearer content display. Plain text is also formatted to highlight important information, as indicated in Example 27, with bold and italic texts.
@@
Error Fixing (61). Error fixing in the domain of README documentation presentation involves addressing two perspectives: grammar and typo fixing, as well as Markdown syntax fixing. Documentation writers might inadvertently make errors when composing README documents, leading to additional time needed for meticulous checking.
@@
Regarding the style of text rephrase and editing, we observed instances where efforts were made to simplify the sentences or render them more succinct. In commit Example 30, the phrase instead of ‘pip install .‘ was deleted. Since there already exists the correct code command, providing an incorrect command is unnecessary and could potentially lead to users’ misuse. On the other hand, there are also light updates that perform less apparent modifications. For example, in commit Example 31, the phrase doesn’t necessarily work is revised to might not work as expected. The modified text included the behaviour that code does not work, highlighting the potential limitations.
@@
4.6 External Resources Management
External Resources refer to other documentation and non-documentation artefacts that are not embedded within the project’s README files, but are accessible only via the provided URL links. In total, 366 of 1,168 data points are related to this category. In this category, we synthesised three concepts, which are discussed as follows:
@@ 4.6 External Resources Management
Project Non-documentation Artefacts Management (136). Apart from external documents, other artefacts (e.g., source code, release package, data, and configuration file) are closely related to installation-related tasks. Typically facilitated by Git, the most frequent modification for non-documentation artefact is on source code access. This access is a crucial step to enable the source code installation method. Release packages, such as the jar package for Java projects, are also actively maintained for users to get access to different versions of the release. Additionally, some projects organise other URLs, such as links to configuration files (e.g., Dockerfile), providing quick access to corresponding files, such as the update in commit Example 34. This practice provides quick access to editable files, which helps with local development.
@@
Summary of external resources management results: Project external documentation management is related to providing links for other detailed project document, which helps avoid lengthy README files. Project non-documentation artefacts refers to URLs for project artefacts including source code, executable and other configuration files. Furthermore, the concept of third-party external resources management corresponds to updates for URLs for external requirements of the software project.
4.7 Cross-Cutting Concerns
@@
(1) Version Management. As described in previous sections, version management is a notable area where substantial efforts are invested when updating README files. With the development of the project, new versions are released to incorporate additional features, which leads to the active update for version information. In summary, we identified that version management primarily focuses on project version, dependency version, and operating system version.
In terms of project version, information is provided for the latest stable version and the developing version. This version information is usually updated within installation-related instruction, project running instructions, URLs for accessing project artefacts, and plain text manifest. Additionally, previous versions are sometimes provided, and outdated versions are included as help information. Regarding dependency version updates, the programming language version will be updated when a newer version is incorporated into the software project, e.g., migrating from Python 2 to Python 3 or adding support for the latest version of JVM. As for other dependencies, although version information is not always displayed in the README file for every repository, their versions are also updated accordingly in the case that they are displayed, to reflect the changes within each project. As indicated in the previous example, operating system version updates reflect project changes that introduce a newer compatible version. In addition, when a new compatible operating system is added to the project, it might be necessary to document specific instructions on different software development tasks.
@@
5 Implications and Discussion

### v3

@@ IV-C Post-installation Instruction
In addition, OSS projects sometimes contain multiple runnable modules, each serving a specific aspect of the project, such as server/worker modules in distributed system projects and training/inference modules in machine learning projects. Properly documenting each component’s execution instruction is essential for users to make use of the software product. Within our analysed data, we identified in Example 15 that additional execution commands were added for other blocks of functions.
@@
Summary of post-installation instruction results: The most dominant concept in this category is project running instruction. Changes in this concept are found to be related to updating parameter options and file paths to the executables. Other concepts are less frequently encountered but they all constitute the installation-related software development tasks.
IV-D Help Information
@@
Tutorial Updates (70). Limited in space and purpose, README files typically do not offer comprehensive usage instructions, such as API documentation. We observed that most updates within this concept involve providing code demos for the basic usage of a project. Most documentation changes are trivial, merely made to align with the source code modifications. In Example 17, updates are made within the code demos, altering all naming paradigms for the methods of using the model.
@@
Summary of help information results: The most frequent updates in this category are the explanation information updates, followed by notes and troubleshooting updates and tutorial updates. Unlike instructional updates, updates in this category focus more on providing clarifications and examples, but are also less frequent.
IV-E Document Presentation
Different from the previous sections, document presentation updates involve less content modification. Instead, the emphasis is on organising the documentation to enhance visual presentation and reading comprehension. In total, 477 of 1,168 data points are associated with this category, indicating a relatively heavy documentation effort. In this category, we synthesised five concepts, which are discussed as follows:
@@ IV-E Document Presentation
Content Formatting and Cleaning (119). README content formatting involves using Markdown syntax to either highlight or organise the content. Various information is commonly formatted, including large code blocks in Example 24, inline code blocks in Example 25, and URLs in Example 26. All of these formatting operations contribute to enhancing the presentation of the documentation, making the README files easier to follow. Specifically, the formatting on code blocks enables better highlighting of code components, while the formatting on URLs facilitates enhanced accessibility and clearer content display. Plain text is also formatted to highlight important information, as indicated in Example 27, with bold and italic texts.
@@
Presentation Fixing (61). Presentation fixing in the domain of README documentation presentation involves addressing two perspectives: grammar and typo fixing, as well as Markdown syntax fixing.
@@
Regarding the style of text rephrasing and editing, we observed instances where efforts were made to simplify the sentences or render them more succinct. In commit Example 30, the phrase instead of ‘pip install .‘ was deleted. Since there already exists the correct code command, providing an incorrect command is unnecessary and could potentially lead to users’ misuse. On the other hand, there are also light updates that perform less apparent modifications. For example, in commit Example 31, the phrase doesn’t necessarily work is revised to might not work as expected. The modified text included the behaviour that the code does not work, highlighting the potential limitations.
@@
IV-F External Resources Management
External Resources refer to other documentation and non-documentation artefacts that are not embedded within the project’s README files but are accessible only via the provided URL links. In total, 366 of 1,168 data points are related to this category. In this category, we synthesised three concepts, which are discussed as follows:
@@ IV-F External Resources Management
Project Non-documentation Artefacts Management (136). Apart from external documents, other artefacts (e.g., source code, release package, data, and configuration file) are closely related to installation-related tasks. Typically facilitated by Git, the most frequent modification for the non-documentation artefact is on source code access. This access is a crucial step to enable the source code installation method. Release packages, such as the jar package for Java projects, are also actively maintained for users to get access to different versions of the release. Additionally, some projects organise other URLs, such as links to configuration files (e.g., Dockerfile), providing quick access to corresponding files, such as the update in commit Example 34. This practice provides quick access to editable files, which helps with local development.
@@
Summary of external resources management results: Project external documentation management is related to providing links for other detailed project documents, which helps avoid lengthy README files. Project non-documentation artefacts refer to URLs for project artefacts including source code, executables and other configuration files. Furthermore, the concept of third-party external resources management corresponds to updates for URLs for external requirements of the software project.
IV-G Cross-Cutting Concerns
@@
(1) Version Management. As described in previous sections, version management is a notable area where substantial efforts are invested when updating README files. With the development of the project, new versions are released to incorporate additional features, which leads to the active update for version information. In summary, we identified that version management primarily focuses on project versions, dependency versions, and operating system versions.
In terms of the project version, information is provided for the latest stable version and the developing version. This version information is usually updated within installation-related instructions, project running instructions, URLs for accessing project artefacts, and plain text manifest. Additionally, previous versions are sometimes provided, and outdated versions are included as help information. Regarding dependency version updates, the programming language version will be updated when a newer version is incorporated into the software project, e.g., migrating from Python 2 to Python 3 or adding support for the latest version of JVM. As for other dependencies, although version information is not always displayed in the README file for every repository, their versions are also updated accordingly in the case that they are displayed, to reflect the changes within each project. As indicated in the previous example, operating system version updates reflect project changes that introduce a newer compatible version. In addition, when a new compatible operating system is added to the project, it might be necessary to document specific instructions on different software development tasks.
@@
IV-H README Update Behaviours
Last, we summarise the overall update behaviours based on the derived codes. The most frequently updated category is ”Pre-installation instruction,” with dependency setup being the most modified concept. Documentation presentation is the second most focused category, where maintainers primarily work on modifying the documentation structure and formatting the documents for better readability. Additionally, updates to instructional commands significantly outnumber updates to help information, with nearly five times as many. These observations indicate that OSS documentation maintainers invest more effort in updating dependencies and documentation structures. They also focus more on providing instructional information rather than help information such as troubleshooting in README files.
Additionally, we distinguish our base codes with the operations of “Addition”, “Modification”, and “Removal”, with the statistics displayed in Table II. Unlike the section-level updates shown in Figure 4, we categorise them based on more detailed modification behaviours. Specifically, whether the modification introduces new information, updates existing information, or removes old information. Seen from the table, apart from document presentation, the largest proportion of update behaviour involves “addition”. This indicates that more information is being incorporated into README files as the software development process progresses, requiring documentation maintainers to spend time assembling new information to reflect system changes.
TABLE III: Quantitative evidence on update categories and information presented
Update Information
Pre-installation Installation Post-installation External resources Help information
Pre-installation instruction 27 19 14 17 11
Installation-related instruction 15 28 17 18 7
Post-installation instruction 19 26 29 10 10
External resources management 20 26 15 26 9
Help information 16 24 20 15 20
Document presentation 16 30 17 13 7
We also provide quantitative evidence on information already present in the first version of a README when different categories of updates occur. Specifically, we sampled 30 documentation updates from each category and annotated the first version of each README to identify the occurring information. As approximately 60% of the data involves only one type of update category, and to better illustrate the distribution of sections across independent update behaviours, we limited our sampling to data containing only a single update category. Table III shows the frequency of each section present in the previous README version (totalling 30 samples) when each follow-up modification occurs, highlighting the relationship between section completeness and modification patterns.
From the table, it is evident that all types of information generally exist in the initial version of a README file. Notably, in the first version across various update behaviour categories, installation information is the most frequently included content, except in cases where pre-installation instructions are updated. Moreover, for each update behaviour category (except for document presentation), related content was already present in the initial version of the README. This finding suggests that practitioners often update documentation incrementally rather than introducing entirely new content.
V Triggers for Documentation Updates
Additionally, we analysed the factors that triggered documentation updates across the entire set of 1,163 documentation updates. Specifically, we followed a series of steps to identify these triggers, stopping once the trigger was identified or reaching the final step. The steps are as follows: (1) read and understand the document update, (2) review the previous README version for obvious issues (e.g., syntax errors, incomplete instructions, presentation issues), (3) examine the commit message to determine whether it is directly linked to a pull request or issue, or if it includes the motivation for the update, (4) if the commit is associated with a pull request/issue, review the first block of the description and initial discussion of that pull request, (5) examine any additional pull requests linked in one level in the discussion thread (non-recursive), (6) check the most recent release for the repository corresponding to this README version, (7) search for modified keywords such as package names and file paths using the GitHub search function, and (8) if none of the above methods yields an answer, we mark the trigger as “no explicit evidence” as we were not able to identify explicit evidence for what triggered the update.
This procedure for identifying triggers was developed based on an initial set of pilot data, discussed and refined until reaching consensus among all authors. The design of this process aims to identify triggers using the most commonly used channels in open-source projects while avoiding spontaneous and endless searching, ensuring that our method is replicable. After that, the first author analysed the rest of the triggers and flagged uncertain entries for further discussion.
Following this procedure, we identified 614 triggers from the entire 1,168 documentation updates, while the rest of the documentation updates are annotated with “no explicit evidence”. We then performed open card sorting [53] to group similar triggers into the same group and assign meaningful names to it. Specifically, we identified 11 different types of triggers. These 11 triggers are further grouped into 3 bigger themes, namely 1) Errors in the previous documentation, 2) Changes in the codebase, and 3) Need for documentation improvement. In the following, we introduce each type of trigger, indicating in parentheses how many instances we identified for each type.
V-A Errors in the previous documentation
Presentation issues in the previous version of README (93). Documentation maintainers update their documentation when identifying issues within the previous version of a README that impact its presentation. For example, the commit [54] added a white space between the hashtag and the section title, as the previous version of the README violated the markdown syntax for displaying headers.
Incomplete contents in the previous version of README (78). In this case, documentation maintainers left the documentation clearly unfinished in the previous version. Documentation debt, such as explicit TODOs [55] or documentation with only skeletons of section headers [56], are strong indicators of unfinished jobs, suggesting that new content should be added in accordance with the context provided. Meanwhile, when the README document is added to the project at the beginning (i.e., the first version of the README), a lot of content is needed, and our proposed template might be a good starting point.
Instruction or content issues in the previous version of README (48). Instruction or content errors severely impact downstream users’ experience in installing and using the project, and developers should quickly identify the existence of such issues and fix them. For example, commit [57] removes command --system-site-packages. The error in this command is reported in issue 1 in this repository by a user and its fix is later reflected in the document.
Outdated elements in previous README (12). Software documentation gets silently outdated during the software development process. When developers realise the existence of such outdated pieces, they make adjustments in the README documents accordingly. For example, commit [58] updates the Python version to 3, while the migration from Python 2 to Python 3 was done in a pull request three years before this documentation update [59], indicating the existence of such an outdated element in the README file.
V-B Changes in the codebase
Update of software version (131). This is the most frequent trigger. When new releases of a software are made, which can usually be observed in the repository release channel or release log, version information in the documentation gets incremented as well, as showcased in this example [60].
Add or modify software functionality (83). This relates to large changes in the codebase, usually accompanied by adding, updating, or deleting multiple files. As a result of such updates, software functionalities such as installation or running instructions might be updated. This requires the documentation authors to carefully account for the changes to be updated in the README. For example, pull request [61] made large updates with various functionalities, and the new OS support is reflected in the README within the commit [62]. Such information is difficult to summarise from the large amount of code changes, while clear pull request descriptions can help developers easily track what should be documented in the README files.
Command updates or file path updates (76). Simple command and file path updates are short code snippet modifications that can be easily tracked to the corresponding files and directories. When such code elements get updated during the development, README changes are also required to keep the documentation up-to-date. For example, commit [63] moves the directory configs from the root folder to the code2seq folder, thus triggering updates to the file path present in the documentation.
Update of software dependencies (52). Dependencies evolve as the software evolves. Updates to dependencies also triggers documentation updates. For example, commit [64] updates the dependency Confluent version to 3.1.2, while such an update can be identified within another commit [65] within the same pull request to update the build.gradle file. In fact, the dependency update information can usually be found within the build file or requirement file in the same repository.
V-C Need for documentation improvement
Self-motivated documentation improvement (50). Some projects inspect their README document periodically to evaluate its quality and propose updates to make it easier to understand. For example, a lot of effort was spent in pull request [66] to make the README file better. Incorporating our proposed template could help mitigate the effort spent on this aspect.
Request of documentation updates by downstream users (13). Sometimes downstream users proactively approach the developers, requesting an explanation or enhancement of the documentation. The GitHub Issue channel is suitable for monitoring such concerns raised by users. For example, issue [67] asked for an installation option on Ubuntu and Debian, and the instruction was added later in commit [68].
VI README Template and its Validation
Figure 10: Proposed README Template
As we discovered that most categories contain more addition operations and that significant efforts are spent on documentation presentation adjustments by practitioners, we created a README template based on the derived taxonomy. This template encompasses all essential installation-related contents, aiming to provide a comprehensive checklist and reduce practitioners’ maintenance efforts for README files. In this section, we describe our proposed template and validate it through a survey conducted on Prolific.
VI-A README Template Content
Figure 10 presented the suggested template. It covers basic instructions including prerequisites, installation steps, usage, external documents, version history, and help and support sections. These sections are linked to the taxonomy derived from our RQ1, and we have explicitly labelled them in the figure. In the prerequisites section, we incorporate programming language version and dependencies for the project. URLs for the official websites of the dependencies are also provided. The installation steps include both source code level installation and published package level installation. The usage section covers basic commands for running the project, as well as some advanced options to display the capability of the software. External documents provide links to external comprehensive documents. The version history section provides the version information of this project, while the help and support section provides most common FAQs and channels for users to interact with. In cases where specific functionalities are yet to be implemented, these placeholders can be seen as a form of self-admitted document debt, serving as reminders for document maintainers to complete the necessary information.
VI-B Template Validation Process
To further verify whether our proposed template is perceived as useful by the documentation users, we conducted a survey on Prolific. We applied our proposed template to the original README documents by performing rewrites. To avoid subjectivity in the rewriting process, we instructed ChatGPT to handle this task. Specifically, we provided the original README file and the template, along with two additional instructions: (1) do not omit any details from the original document, and (2) do not add any information not present in the original document. We used ChatGPT 4o throughout this procedure. Only one entry of instruction is provided to ChatGPT to collect the generation, without further tuning or additional feedback.
For the evaluated samples, we first randomly selected 20 repositories. From each repository, we randomly chose one commit to ensure each document was selected only once per repository. We then instructed ChatGPT to apply our proposed template to these documents. We refer to the original documents as “original” documents and the generated documents as ”augmented” documents. In total, we obtained 20 “original” to “augmented” document pairs for evaluation.
Table IV shows the number of projects containing the top two levels of headers in both the “original” and “augmented” documentation. For these two most frequent headers, the majority of original documents feature these top two header levels, indicating that the original README versions are already structured. The augmented version, however, contains slightly fewer top-level headers. As a result, the evaluation is not biased by having more structure in the augmented version, allowing for a focus on the content and organisation of our proposed template.
TABLE IV: Counts for Header Scales for Sampled Repositories
Header Scale $\langle h1\rangle$ $\langle h2\rangle$
Original Version Count 18 16
Augmented Version Count 13 15
We conducted a Prolific survey by dividing the 20 document pairs into 10 different surveys, with each survey taken by three different participants. We performed a sequential survey release as in [35] to prevent the same individual from participating multiple times. Meanwhile, we followed a study design similar to Nadi and Treude [69] by inserting a “quality gate” in the survey. After reading the documents, we asked the participants to answer a simple question, e.g., “Is Gradle used in the above-mentioned project?”. We filtered out participants who did not answer the question correctly.
To evaluate how users perceive the documents, we adopted an evaluation framework proposed by Treude et al. [13], which provides a ten-dimensional framework for assessing software documentation quality. Table V displays these dimensions as well as questions in this paper. At the beginning of the survey, we first provided the definitions of these ten dimensions and asked the participants to read through them. After that, they were requested to read through the original and augmented documents and provide scores on these ten dimensions. To prevent biased judgements, the documents were anonymised regarding their status as original or augmented. Likert scale is used to mark each of the aspects, with a score of 1 for strongly disagreeing and a score of 5 for strongly agreeing. Additionally, we revealed the template at the end of the survey and with an open-ended question aiming to gather participants’ opinions about it. We added a sample as well as all README file pairs in our replication package.
TABLE V: Dimensions of software documentation quality
Dimension Question
Quality How well-written is this document (e.g., spelling, grammar)
Appeal How interesting is it?
Readability How easy was it to read?
Understandability How easy was it to understand?
Structure How well-structured is the document?
Cohesion How well does the text fit together?
Conciseness How succinct is the information provided?
Effectiveness Does the document make effective use of technical vocabulary?
Clarity Does the document contain ambiguity?
Figure 11: Survey Result
VI-C Demographics of Annotators
We first applied predefined filters on Prolific, requiring participants to be in the Information Technology (IT) employment sector and be proficient in English. We asked the same question about their employment again at the beginning of our survey and filtered out 18 participants who did not answer this question consistently with their Prolific registration. Our “quality gate” question further filtered out four participants.
In the survey, we asked participants about their job roles and how many years they have worked in the IT field. For the 30 participants who passed all filters and submitted their responses, they have on average 6.9 years of experience working in IT, with a maximum of twenty years and a minimum of one year. Figure 12 shows the demographic distribution, with the largest proportion of participants being “developers”, and the legend sorted in descending order based on the number of participants. We computed the inter-rater agreement among the evaluators, which reached 0.34 for Krippendorff’s alpha [70] coefficient overall. This indicates that the evaluators reached moderate agreement, while the evaluation remains somewhat subjective.
Figure 12: Job Roles of Annotators
VI-D Template Validation Result
Seen from Figure 11, the average scores for the “augmented” document are consistently higher than those of the “original” documents. The figure is sorted based on the difference between the average scores of these two types of documents. The most improved dimensions are structure, understandability and conciseness, with a gain of 0.80, 0.53, and 0.53 of the Likert average score, respectively. Meanwhile, the increases in clarity and consistency are less noticeable.
We also performed Wilcoxon signed-rank tests [71] on all dimensions to determine whether the differences are statistically significant. As indicated in the figure, eight out of ten dimensions are with a p-value below 0.05, indicating statistically significant differences. We further computed the Cohen’s d effect size [47] for the ten evaluation dimensions. The results indicate that the adoption of our proposed template significantly enhances the perceived structure with a large effect, while significantly enhances understandability, conciseness, readability, quality, cohesion, appeal, and effectiveness with medium effect. The improvements in clarity and consistency are small.
Therefore, based on the quantitative results of our survey, we find that applying our proposed template increases the quality of the generated documents, particularly in terms of structure, understandability, and conciseness, without compromising clarity and consistency, as perceived by document readers.
VI-E Practitioners’ Perspectives from Pull Requests
To further assess the proposed README template and methodology, we conducted an additional study to gather developers’ opinions. Specifically, we applied the proposed template to the latest version of README files and submitted pull requests to the respective repositories, requesting the suggested changes. Out of the 400 repositories initially considered, we excluded those without any commits in the past year, narrowing the sample to 219 active repositories. From these, we randomly selected 30 repositories to conduct the experiment.
In the pull request, we included a commit with the updated README file that integrates our proposed template. Along with this, we provided a description of the potential benefits of adopting the template and included a screenshot of the raw template for reference. At the time of writing, among eight repositories that responded and were comfortable with LLM-generated improvements to their README files, six projects responded positively, while two projects responded negatively. Four repositories expressed reluctance to use LLM-generated content for documentation or questioned the purpose of unsolicited pull requests.
Positive feedback. Figure 13 shows the merged pull request to project aerleon/aerleon. [^6] https://github.com/aerleon/aerleon/pull/378 Similarly, project spring/uberserver and actinia-org/arctinia-core also directly merged our proposed change without further comment. [^7] https://github.com/spring/uberserver/pull/400 Project apache/openmeetings responded positively with “I really like the new look of the readme” and suggested some customised changes. [^8] https://github.com/apache/openmeetings/pull/196 The pull request was then merged into the codebase after we made the suggested changes. Additionally, project actris-cloudnet/cloudnetpy adopted part of our proposed changes while leaving other parts of their README unchanged. [^9] https://github.com/actris-cloudnet/cloudnetpy/pull/110 Specifically, they adjusted section header hierarchies to improve consistency. For project vespa-engine/pyvespa, [^10] https://github.com/vespa-engine/pyvespa/pull/976 the maintainer commented, “I think we can add some of the suggestions, but some also are not correct/ helpful.”, and further suggested that “do a human review before creating a PR”. This feedback highlights the potential for hallucinations in ChatGPT-generated suggestions and underscores the importance of human validation to ensure correctness when updating documentation.
Figure 13: Example of Merged Pull Request
Negative feedback. Out of the two negative responses, one reason for rejection is the lack of project-specific details in the updated version. For example, project 3dcitydb/importer-exporter mentioned that “Some sections that are important for us, like the partners and supporters and additional information about the technical standards behind the project are missing”. [^11] https://github.com/3dcitydb/importer-exporter/pull/312 Additionally, both repositories viewed the proposed updates as offering insignificant improvement.
Therefore, we see that our proposed method can be adopted by open-source repositories and is endorsed by several merges and positive comments from real-life projects. However, the practitioners also raised critical points regarding the completeness and correctness of the generated documentation. Semi-automated documentation updates with humans in the loop might be a solution for such concerns while reducing manual efforts to a large extent.
It is important to note that developing a directly applicable pipeline for applying the template to README files is not the primary focus of this paper. Instead, the validations obtained through user feedback and this pull request study serve as a proof of concept, demonstrating that the template has the potential to make real-world impact. Future research should focus on further refining the template and developing a comprehensive methodology for its application, ensuring it can be seamlessly integrated into real-world production workflows.
VII Implications and Discussion